• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 16:00:34 on December14,2008

    Entry number 253399

    keyword=Day Shift Summary

    SL: Diana Parno
    TO: Xiaoyan Deng
    3rd: Chunhua Chen

    Shift began with production running smoothly for \pi+V.

    We experienced a problem for the first several runs whereby the pop-up windows did not appear for run starts and stops. This was later solved with Kalyan's help by killing coda, re-logging into the a-onl account, and starting coda again.

    Production runs:
    4477 (in progress from last shift) ended with 0.04 C (estimated)
    4478 0.04 C (estimated)
    4479 (ended early to restart CODA)
    4480 (very short -- 2 min)
    4481 0.0405 C (CODA working again)
    4482 0.04002
    4483 0.03997

    --- 4484: RICH Pedestal run

    4485 0.04 (estimated)
    4486 0.04034
    4487 begun; will finish during swing shift

    Shift ended in smooth production running.

    Total beam charge is estimated at 0.28 C.